Question: Consider a bag containing 5 green, 7 black and 4 white marbles. If three marbles are drawn randomly then what is the probability that none of them is white?
Answer #1

Number of white marbles = 4

Total number of marbles = 5+7+4 = 16

Number of non white marbles = 16 - 4 = 12

Number of ways to draw r items from n given items is given by, nCr = n!/(r! x (n-r)!)

P(none of the three marbles drawn are white) = Number of ways in which 3 non white marbles can be drawn / Number of ways in which total 3 marbles from be drawn

= 12C3 / 16C3

= 220/560

= 11/28

= 0.3929